### v1

# Vector interaction bounds in NJL-like models from LQCD estimated curvature of the chiral crossover line

Abstract
We obtain improved bounds on both the flavor-independent and -dependent vector interactions in a $2+1$-flavor Nambu–Jona-Lasinio (NJL) model using the latest precise LQCD results of the curvature coefficients of the chiral crossover line. We find that these lattice estimated curvature coefficients allow for both attractive and repulsive types of interactions in both the cases. With this constrained ranges of vector interactions, we further predict the behavior of the second $(\kappa_{2}^{B})$ and fourth $(\kappa_{4}^{B})$ order curvature coefficients as a function of the strangeness chemical potential $(\mu_{S})$. We observe that the flavor mixing effects, arising from the flavor-independent vector interaction as well as from the ’t Hooft interaction, play an important role in $k_{2}^{B}$. We propose that the mixing effects due to the vector interaction can be separated from those arising from the ’t Hooft interaction by analyzing the behavior of $k_{2}^{B}$ as a function of $\mu_{S}$. Finally, we locate the critical endpoint in the $T-\mu_{B}$ plane using the model-estimated ranges of vector interactions and find the model’s predictions to be consistent with the latest LQCD bounds.
I Introduction
Over the last few decades, substantial progress has been made in understanding the strong interaction system at low energies, where the spontaneous breaking of (approximate) chiral symmetry plays a crucial role. This broken symmetry is restored as the temperature $(T)$ and/or baryon chemical potential $(\mu_{B})$ of the system increases. The transition from the broken to the restored symmetry phase is crucial for understanding phenomena such as the quark-gluon plasma (QGP) produced in relativistic heavy ion collisions (HICs), as well as the early universe in the context of the Big Bang and compact astrophysical objects like neutron stars, which may have a quark core [1].
First-principle lattice QCD (LQCD) calculations have made significant progress over the past two decades, greatly enhancing our understanding of the chiral transition at zero or low baryon chemical potential $(\mu_{B})$. However, the lattice artifacts arising from the well-known sign problem restrict the extension of such studies at large $\mu_{B}$. In the current scenario, techniques such as Taylor expansion around $\mu_{B}=0$ [2, 3, 4] or analytical continuation from imaginary $\mu_{B}$ [5, 6] have produced reliable results for $\mu_{B}/T\leq 3$ [7, 8]. Additionally, functional methods like the Dyson-Schwinger equations (DSE) [9] and functional renormalization group (FRG) [10] offer trustworthy results for higher $\mu_{B}$ values through their improved truncation schemes.
On the other hand, effective models constructed based on the underlying symmetries of the theory exhibit broad applicability across the $T-\mu_{B}$ plane. In scenarios where reliable first-principle calculations are lacking, these models can provide significant insights, specifically into the low-energy sector of QCD. Even with advancements in first-principle results, these models continue to be relevant, particularly for exploring extreme conditions such as those characterized by high baryon density [11].
The Nambu–Jona-Lasinio (NJL) model serves as a quintessential example of an effective model, extensively used to investigate the breaking and restoration of chiral symmetry, as well as the influence of external parameters on the underlying dynamics [12, 13]. Initially proposed by Y. Nambu and G. Jona-Lasinio with nucleon degrees of freedom [14, 15], the model was later formulated in terms of quark degrees of freedom following the discovery of quarks [16, 17]. As a low-energy model, one can consider all possible interactions up to a given order, constrained by the underlying symmetries. Even with all the possible terms within the mean-field approximation, the low-energy dynamics can still be captured using only a few parameters, as demonstrated in Ref. [18].
Among the various types of interactions, vector interaction plays a crucial role. Most importantly, in a dense medium, it becomes essential to account for a non-zero vector interaction due to its direct coupling with the number density operator. As a result, even if we start with a zero strength of the vector interaction, the non-zero density environment induces a finite strength for this interaction. As it becomes relevant only at finite density, it cannot be determined using vector meson properties in the vacuum, and its strength remains an uncertain quantity. This contrasts with other interactions, such as the scalar interaction, which can be fixed based on the vacuum properties. Consequently, in much of the existing literature, the strength of the vector interaction ($G_{V}$) is varied in units of the scalar interaction strength $(G_{S})$. It is also worth noting that the sign of $G_{V}$ is not universally agreed upon and can be either positive or negative, leading to repulsive or attractive interactions, respectively [19].
The vector interaction has significant effects on the restoration of chiral symmetry at non-zero baryon chemical potential ($\mu_{B}$) and directly influences the location of the critical endpoint (CEP) in the $T-\mu_{B}$ plane. Furthermore, the curvature coefficients of the chiral crossover line around zero $\mu_{B}$ are strongly affected by the vector interaction. These features have been exploited to understand the role of vector interaction in QCD and to constrain its range in effective model frameworks. For instance, Refs. [20, 21, 22, 23, 24] used different effective model frameworks, including variants of the NJL model, to estimate the vector interaction strength primarily from the LQCD-derived curvature of the chiral crossover line, among other observables. Due to the lack of precise estimates from LQCD, the bounds remain broad but predominantly suggest a repulsive interaction, with the chiral transition becoming a crossover throughout the phase diagram at a sufficiently strong positive interaction strength [19]. Vector interactions also have a profound impact on bulk thermodynamic properties at high densities, contributing to an acceptable equation of state (EoS) for astrophysical objects, such as neutron stars [25] or hybrid stars [26].
In this article, we employ a $2+1$-flavor NJL model to revisit the problem of determining the strength of the vector interaction, utilizing the improved and more precise LQCD data on the curvature [4]. Our motivation is drawn from a recent study [27], which found that the curvature coefficients of the crossover line obtained within a $2+1$-flavor NJL model align remarkably well with LQCD simulations and exhibit minimal dependence on model parameters. Additionally, it was proposed that determining the second-order curvature coefficient of the $T-\mu_{B}$ line ($\kappa_{2}^{B}$) as a function of the strangeness chemical potential ($\mu_{S}$) on the lattice would enhance the predictability of NJL-type models. We extend this framework by introducing both flavor-independent and flavor-dependent vector interactions, with coupling strengths $G_{V}$ and $g_{V}$, respectively.
We varied the vector interaction strengths $G_{V}$ and $g_{V}$ in units of $G_{S}$ across positive and negative ranges and evaluated the corresponding $\kappa_{2}^{B}$ and $\kappa_{4}^{B}$. Using precise lattice QCD data for $\kappa_{2}^{B}$, we narrowed down the range of these vector interaction strengths. However, $\kappa_{4}^{B}$ estimates from LQCD are consistently zero within uncertainties, offering limited guidance. Lattice results that impose the strangeness neutrality condition (zero net strange quark) provide an additional method for constraining the data with a finite strangeness chemical potential, $\mu_{S}$. This motivates further exploration of the impact of $\mu_{S}$ on the curvature coefficients $\kappa_{2,4}^{B}$ in the presence of vector interactions, previously studied in Ref. [27] without them. This analysis enhances our understanding of flavor mixing between the light and strange quark sectors. In the NJL model, the ’t Hooft determinant term generally captures flavor mixing, while a flavor-independent vector interaction couples the two sectors via their number densities as well. By examining the variation of $\kappa_{2}^{B}$ with $\mu_{S}$, we propose a novel approach to distinguish the contributions from these two mechanisms of flavor mixing and outline potential implications. With the narrowed range of $G_{V}$ and $g_{V}$, we finally provide the location of the critical endpoint (CEP) in our model, which is consistent with the existing LQCD bound [7, 8], however, differing significantly from functional methods’ present predictions [28, 29].
The article is presented in the following way. In section II, we provide the details of the NJL model as required for the present study and briefly discuss the procedure to estimate the curvature. In section III, we show the results and discuss the estimation of the curvature coefficients in different physical conditions. Finally, in section III.4, we find the locations of the CEP within the allowed range of parameters before we conclude in section IV.
II Formalism
II.1 Details of the model
We start with the Lagrangian density of a $2+1$-flavor NJL model, including the vector interaction, which can be written as [30, 19, 31]
$\mathcal{L}$ $=\bar{q}\left(i\not{\partial}-\hat{m}+\gamma^{0}\hat{\mu}\right)q+{G_{S}}\sum_{a=0}^{8}\left[(\bar{q}\lambda_{a}q)^{2}+(\bar{q}i\gamma^{5}\lambda_{a}q)^{2}\right]$
$-8K[\det(\bar{q}P_{R}q)+\det(\bar{q}P_{L}q)]$ (1)
$-\begin{cases}G_{V}(\bar{q}\gamma^{\mu}q)^{2}\\ g_{V}\sum_{a=0}^{8}\left[(\overline{q}\gamma^{\mu}\lambda_{a}q)^{2}+(\overline{q}i\gamma^{\mu}\gamma^{5}\lambda_{a}q)^{2}\right]\ \end{cases}$
where the quarks are represented by $q$ as $q^{\text{T}}=(u,d,s)$. $\hat{m}=\text{diag}(m_{u},m_{d},m_{s})$ and $\hat{\mu}=\text{diag}(\mu_{u},\mu_{d},\mu_{s})$ are the mass and chemical potential matrices in the flavor space, respectively. $G_{S}$ is the strength of the Lorentz scalar interaction that is invariant under global $U(3)\times U(3)$ symmetry, where $\lambda_{a}$’s are the Gell-Mann matrices, the generator of $U(3)$ symmetry in the fundamental representation with normalization $\text{Tr}[\lambda_{a}\lambda_{b}]=2\delta_{ab}$ and $\lambda_{0}\propto\mathbb{1}_{3\times 3}$. To take into account the effect of axial anomaly, one explicitly breaks the $U(1)_{A}$ symmetry by introducing the Kobayashi–Maskawa–’t Hooft (KMT) interaction [32, 33] with coupling $K$, where $P_{L/R}=(1\mp\gamma_{5})/2$ is the chiral projection operator. We have further considered two types of vector interaction—  flavor independent and flavor dependent [30]. In that regard, they are two different models, and we call them Model-I (flavor independent, $G_{V}$) and Model-II (flavor dependent, $g_{V}$). Next, we obtain the bosonized action by introducing the meanfields (MFs).
From the meanfield Lagrangian, one can obtain the grand canonical potential evaluating the partition function. To introduce temperature, we have used the Matsubara formalism [34, 35, 36], which connects temperature to the zeroth component of the four-momentum. The free energy can be written as
$\Omega(T,\mu)=\Omega_{\text{cond}}+\Omega_{\text{vac}}+\Omega_{\text{med}},$ (2)
where $\Omega_{\text{cond}}$ represents the condensation energy, $\Omega_{\text{vac}}$, the vacuum energy or zero-point energy and $\Omega_{\text{med}}$ denotes the medium contribution. The explicit expressions are
$\Omega_{\text{cond}}$ $=2G_{S}\sum_{i}\sigma_{i}^{2}-4K\prod_{i}\sigma_{i}-\begin{cases}G_{V}\left(\sum_{i}n_{i}\right)^{2}\\ g_{V}\sum_{i}n_{i}^{2}\end{cases},$
$\Omega_{\text{vac}}$ $=-2N_{c}\sum_{i}\int_{\Lambda}\frac{d^{3}p}{(2\pi)^{3}}E_{i}(p)\;\;\;\mathrm{and}$
[…]
$\left.+\ln\quantity[1+e^{-(E_{i}(p)-\tilde{\mu}_{i})/T}]\right).$ (3)
with $N_{c}=3$, the number of colors; $E_{i}=\sqrt{p^{2}+M_{i}^{2}}$, $\sigma_{i}=\expectationvalue*{\bar{q}_{i}q_{i}}$ and $n_{i}=\expectationvalue*{{q}_{i}^{\dagger}q_{i}}$ are the energy, quark condensate and number density of the $i$-th quark, respectively. To render finite contribution from the vacuum term ($\Omega_{\text{vac}}$), we use $3$-momentum cutoff, $\Lambda$, which also represents the scale of the theory. For $i\neq j\neq k\in\{u,d,s\}$, the $i-$th quark effective mass, $M_{i}$, is given by the gap equation
$M_{i}=m_{i}-4G_{S}\sigma_{i}-2K\sigma_{j}\sigma_{k},$ (4)
and the effective chemical potentials are given by
$\tilde{\mu}_{i}=\begin{cases}\mu_{i}-2G_{V}\sum_{j}n_{j}\\ \mu_{i}-2g_{V}n_{i}\,.\end{cases}.$ (5)
One should immediately note the mixing among different flavors through the effective chemical potential in Model-I, which is absent in Model-II. Consequently, in Model-II, the quark chemical potential $(\tilde{\mu}_{i})$ is modified only by the respective number density $(n_{i})$.
To solve for the mean fields, we need to minimize the free energy by solving the gap equations simultaneously
$\frac{\partial\Omega}{\partial\sigma_{u}}=\frac{\partial\Omega}{\partial\sigma_{d}}=\frac{\partial\Omega}{\partial\sigma_{s}}=0.$ (6)
In the present study, we have considered an exact isospin symmetry (ignoring the electric charges of individual flavors), which implies that $\sigma_{u}=\sigma_{d}=\sigma_{l}$. Hence, the individual quark chemical potentials can be expressed in terms of the baryon $(\mu_{B})$ and strangeness $(\mu_{S})$ chemical potentials as
$\begin{split}\mu_{u}&=\mu_{d}=\frac{1}{3}\,\mu_{B}\,,\\ \mu_{s}&=\frac{1}{3}\,\mu_{B}-\mu_{S}\,,\end{split}$ (7)
where, $\mu_{u}$, $\mu_{d}$ and $\mu_{s}$ are chemical potentials for $u$, $d$ and $s$ quarks, respectively.
 $m_{l}$(MeV)  $m_{s}$(MeV)  $\Lambda$ (MeV)  $G_{S}\Lambda^{2}$  $K\Lambda^{5}$
Set I 5.5 135.7 631.4 1.835 9.29
Set II 5.5 140.7 602.3 1.835 12.36
Table 1: Parametrization used in the present work are from Ref. [16] (Set I) and Ref. [17] (Set II).
II.2 Curvature calculation
The crossover line for the $T-\mu_{X}$ plane can be expressed with the following ansatz for smaller values of $\mu_{X}$  [37, 38, 4]:
$\frac{T_{pc}(\mu_{X})}{T_{pc}(0)}=1-\kappa_{2}^{X}\left(\frac{\mu_{X}}{T_{pc}(0)}\right)^{2}-\kappa_{4}^{X}\left(\frac{\mu_{X}}{T_{pc}(0)}\right)^{4}~{}.$ (8)
Here, $X$ denotes conserved charges like baryon charge $B$, electric charge $Q$, and strangeness $S$. The curvature coefficients $\kappa_{2}^{B}$ and $\kappa_{4}^{B}$ in the $T-\mu_{B}$ plane have been explored in lattice QCD, with Taylor expansion method [4] and by analytically continuing imaginary chemical potential results to the real axis [37, 5, 6]. There is very good agreement between the results obtained using these two techniques. There have also been results for the same within other model frameworks such as the NJL model [18], PNJL model [20, 21, 22, 23, 24], perturbative QCD [39], hadron resonance gas (HRG) model [40, 41] and quark-meson model [42, 43, 44, 45, 46, 47].
III Results
Figure 1: The curvature coefficient ($\kappa_{2}^{B}$) as a function of the strengths of vector interactions. Open squares (lines) are obtained using Parameter Set I (Set II), respectively. The bands represent the corresponding Lattice QCD estimations from Ref. [4].
III.1 Constraining $G_{V}$ and $g_{V}$ from $\kappa_{2}^{B}$
Let us first consider the curvature coefficients of the crossover line evaluated with the inclusion of vector interactions. This analysis aids in constraining the interaction strengths $G_{V}$ and $g_{V}$ by comparing them with the lattice QCD estimates of $\kappa_{2}^{B}$ and $\kappa_{4}^{B}$, which are available for the crossover line at small values of $\mu_{B}$. In the NJL model, the pseudocritical temperature ($T_{pc}$) for a given set of $\mu_{B}$ and $\mu_{S}$ values is determined as the inflection point of the order parameter, the light quark condensate, $\sigma_{l}$. We extract the curvature coefficients $\kappa_{2}^{B}$ and $\kappa_{4}^{B}$ by parameterizing the respective crossover lines using the ansatz given in Eq. (8), within the range $\mu_{B}/T_{pc}(0)\leq 1.0$.
 $\kappa_{2}^{B,\mu_{S}=0}$  $\kappa_{2}^{S,\mu_{B}=0}$ $\kappa_{2}^{B,n_{s}=0}$
NJL, Set I 0.0163 0.0134 0.0148
NJL, Set II 0.0162 0.0172 0.0143  [^1] The numerical value mentioned in Ref.[27] is different as that was calculated with less accuracy.
Lattice QCD [4] 0.016(6) 0.017(5) 0.012(4)
Table 2: Estimations of $\kappa_{2}$ for the two parameter sets. The lattice QCD results are taken from Ref. [4].
Let us first summarize the results obtained without the vector interactions in Table 2. We would particularly like to draw attention to $\kappa_{2}^{B,\mu_{S}=0}$ and $\kappa_{2}^{B,n_{s}=0}$. These two curvature coefficients are almost independent of the model parameter sets and have very good agreement with the lattice estimations. On the other hand, $\kappa_{2}^{S}(\mu_{B}=0)$ is sensitive to the choice of model parameters, especially the ’t Hooft determinant, $K$. Due to this robust nature of $\kappa_{2}^{B,\mu_{S}=0}$ and $\kappa_{2}^{B,n_{s}=0}$, we will use them to constrain the strength of the vector interactions and will study the effect of the allowed ranges on $\kappa_{2}^{S,\mu_{B}=0}$. The LQCD data, with which the ranges of $G_{V}$ or $g_{V}$ are constrained, have improved significantly in the last few years. In light of the latest LQCD data, we vary $G_{V}$ and $g_{V}$ in units of $G_{S}$, respectively, and examine the corresponding variations in $\kappa_{2}^{B}$ and $\kappa_{4}^{B}$. In Fig. 1, we illustrate the dependence of $\kappa_{2}^{B}$ on $G_{V}/G_{S}$ and $g_{V}/G_{S}$ for models I and II, respectively.
It is important to emphasize that a negative value of the vector coupling increases the effective chemical potential of the $i$th flavor ($\tilde{\mu}_{i}$) [see Eq. (5)], leading to an increase in the corresponding number density, implying the interaction to be attractive in nature. Similarly, a positive value of the same implies a repulsive-type interaction. Needless to say, the phase diagram in the $T-\tilde{\mu}_{i}$ plane is independent of the strength of the vector interaction. Hence, for attractive interaction, the bare chemical potential ($\mu_{i}$) is less than the effective chemical potential ($\tilde{\mu}_{i}$). This results in a faster decrease of $T_{pc}$ in the $T-\mu_{i}$ plane with increasing $\mu_{i}$ leading to an increase in $\kappa_{2}^{i}$. Conversely, for repulsive interaction, $\mu_{i}$ is greater than $\tilde{\mu}_{i}$ resulting in slower decrease of $T_{pc}$ leading to a decrease in $\kappa_{2}^{i}$. This shift in chemical potential, $\mu_{i}$, accounts for the observed monotonic decrease in $\kappa_{2}^{i}$ as the repulsive strength increases. Following Eq. (7) one can relate $\kappa_{2}^{B}$ of the $T-\mu_{B}$ plane with $\kappa_{2}^{i}$ of the $T-\mu_{i}$ plane by a factor of $9$.
 $\kappa_{2}^{B,\mu_{S}=0}$ $\kappa_{2}^{B,n_{s}=0}$
$G_{V}/G_{S}$ $\left[-0.117,0.212\right]$ $\left[-0.038,0.334\right]$
$g_{V}/G_{S}$ $\left[-0.296,0.552\right]$ $\left[-0.076,0.667\right]$
Table 3: Range of $G_{V}/G_{S}$ and $g_{V}/G_{S}$ constrained from the variation of $\kappa_{2}^{B}$ under different conditions with parameter Set-I.
With the aforementioned observations, we have further constrained the ranges of $G_{V}/G_{S}$ and $g_{V}/G_{S}$ for models I and II using lattice QCD data. In Fig. 1, we have plotted the lattice QCD results for $\kappa_{2}^{B}$, obtained from the HotQCD Collaboration [4]. While similar studies have been conducted by the WB collaboration [6], the HotQCD results are particularly relevant, as they provide data for both the $\mu_{S}=0$ case and the strangeness neutrality condition ($n_{s}=0$). These results are represented by the blue and coral bands, respectively, in Fig. 1.
The points where our estimates intersect with the LQCD bands determine the allowed upper and lower bounds of the vector interaction strengths. The resulting ranges of $G_{V}/G_{S}$ and $g_{V}/G_{S}$ are summarized in Table 3. To understand the effect of model parameters on the estimated range of vector coupling strength, we have used two distinct parameter sets as given in Table 1. From the upper panel of Fig. 1, it is evident that the allowed range of the vector coupling strength is very close to each other for the two parameter sets. It appears that such models exhibit some characteristic features in this regard and are largely independent of the parameters involved [48]. Hence, in the following discussions, we will be presenting all the findings only for parameter set I.
Although Model-I and Model-II are two different models, in certain limits, one can establish a connection between $G_{V}$ and $g_{V}$. For example, with $\mu_{S}=0$, one expects $g_{V}=3.0G_{V}$, assuming degenerate masses for the light and strange quarks leading to $n_{u}=n_{d}=n_{s}$. However, in the scenario considered here, $g_{V}$ is less than $3G_{V}$ due to the much higher mass of the strange quark compared to the light quarks. In contrast, under the strangeness neutrality condition ($n_{s}=0$), the relation becomes $g_{V}=2G_{V}$ as isospin symmetry ensures $n_{u}=n_{d}$. The coupling strength ranges derived in Table 3 follow these conditions.
For small $\mu_{B}/T$ limit, the parametric form of Eq. (8) has a weaker dependence on $\kappa_{4}^{B}$. The estimations from Lattice QCD are zero within the variances, and we have also found a similar trend from our effective model analysis. For a wider range of coupling strength, the $\kappa_{4}^{B}$ lies near zero for both models, providing no additional improvement over the constraint from $\kappa_{2}^{B}$. Such variations have been presented in Appendix A.
We wish to reiterate that in Model I, there is a noticeable crossing between the cases of $\mu_{S}=0$ and $n_{s}=0$ for both parameter sets, whereas such a feature is absent in Model II, where an almost parallel trend seems to exist. In Ref. [27], the authors demonstrated that a finite value of $\mu_{S}$ tends to decrease $\kappa_{2}^{B}$, which is primarily responsible for the lower value of $\kappa_{2}^{B}$ when considering the $n_{s}=0$ case. According to the analytical expression for the effective chemical potential in Model II, as given in Eq. (7), the strangeness neutrality condition explicitly requires $\mu_{S}=\mu_{B}/3$. This introduces a positive shift in $\mu_{S}$ relative to the $\mu_{S}=0$ case, leading to the observed parallel behavior and the absence of any crossing within the explored range of the coupling strength $g_{V}$. In contrast, for Model I, the $n_{s}=0$ condition does not fix $\mu_{S}$ to a specific value but instead makes it dependent on the light quark number densities as $\mu_{S}=\mu_{B}/3-2G_{V}(n_{u}+n_{d})$. This structure implies that at a certain positive $G_{V}$, $\mu_{S}$ will reach zero, explaining the crossing with the $\mu_{S}=0$ case in Model I.
As pointed out in Table 2, the curvature coefficient $\kappa_{2}^{S,\mu_{B}=0}$ is strongly influenced by the model parameters, especially by the ’t Hooft determinant term, $K$. In the absence of a vector interaction, the effect of the strange quark sector propagates to the light quark sector through the mixing term, proportional to $K$ as given in Eq. (4). Without the ’t Hooft interaction, the pseudo-critical temperature $T_{pc}$ would be independent of $\mu_{S}$ resulting in $\kappa_{2}^{S,\mu_{B}=0}=0$. Even though the light quark chemical potential gets modified through $n_{s}$ at finite $\mu_{S}$, the major contributions arise through the flavor mixing in the constituent masses. As a result, we observe that $k_{2}^{S,\mu_{B}=0}$ has a weak dependence on the vector interaction, which is further influenced by the choice of the parameter set and thus cannot be used to obtain any further improvement on the bounds found from $k_{2}^{B,\mu_{S}=0}$ and $k_{2}^{B,n_{s}=0}$. For better visual understanding, we have provided the nature of $\kappa_{2}^{S}$ as a function of $G_{V}/g_{V}$ in Fig. 7 of the Appendix.
III.2 $\mu_{S}$ dependence of $\kappa_{2}^{B}$
Figure 2: The curvature coefficient ($\kappa_{2}^{B}$) as a function of strangeness chemical potential ($\mu_{S}$). $[G_{V}^{-},G_{V}^{+}]$ and $[g_{V}^{-},g_{V}^{+}]$ are the allowed ranges of vector interaction found using LQCD data. Red data points are without the vector interaction i.e., $G_{V}=g_{V}=0$. Black point represents the LQCD estimations of the $\kappa^{B}_{2}$ at $\mu_{S}=0$ [4].
In a recent work [27], some of us investigated the dependence of the curvature coefficients on $\mu_{S}$ and proposed a novel approach for determining the strength of the ’t Hooft determinant interaction, $K$. The ’t Hooft term induces mixing between light and heavy quark flavors, making the precise determination of its strength critical for estimating the degree of flavor mixing in a $2+1$ NJL model. On the other hand, in a $2$-flavor NJL model, this effect gets reflected in isospin-sensitive observables only in the isospin asymmetric scenarios, as shown in Ref. [49, 50]. In the presence of a flavor-independent vector interaction, the effective chemical potential, $\tilde{\mu}_{i}$, depends on the densities of all quark flavors. As a result, the term $G_{V}\sum_{i}n_{i}$ introduces a form of flavor mixing along with the ’t Hooft interaction induced mixing through the constituent mass [see Eq. (4)].
To quantify this, we analyzed the variation of $\kappa_{2}^{B}$ as a function of $\mu_{S}$. The results are presented in Fig. 2. The general trend shows that as $\mu_{S}$ increases, $\kappa_{2}^{B}$ decreases and eventually becomes negative at higher values of $\mu_{S}$ commensurate with the observations of Ref. [27]. We provide estimates constrained by the vector interaction strengths for both Model-I (upper panel, Fig. 2) and Model-II (lower panel), represented by the cyan bands.
One notable feature observed is the variation of $\kappa_{2}^{B}$ with increasing $\mu_{S}$, where the upper and lower curves, corresponding to lowest and highest allowed values of $G_{V}/g_{V}$, respectively, cross each other, intersecting the $G_{V}=0$ line around the same region. In Model-I, this crossing can be interpreted as follows: at the crossing point, the effect of $G_{V}$ becomes negligible. For positive and sufficiently large $\mu_{S}$ compared to $\mu_{B}/3$, the net strange quark density can become negative, potentially offsetting the net light quark density. With a small total quark number density (very close to zero), at low $\mu_{B}$’s, the light quark chemical potential gets a very small modification due to vector interaction, as can be seen from Eq. (5). At this point, the results should coincide with those for $G_{V}=0$, leading to the observed crossing of the two curves (magenta and blue).
Figure 3: Curvature coefficient, $\kappa_{2}^{B}$ as a function of $\mu_{S}$ with $K=0$ for both Model-I and Model-II.
At this juncture, we want to separately assess the impact of flavor mixing. In Ref. [27], it has been observed that this decreasing trend of $\kappa_{2}^{B}$ is due to the flavor mixing between the light and strange quark sectors. The flavor mixing primarily happens for the ’t Hooft interaction term, represented by the $K$, whereas the presence of the vector interaction also provides scope for the flavor mixing, especially in the model I, where the vector interaction is flavor-independent. To separate out the effect arising only from $G_{V}$, we have estimated the $\kappa_{2}^{B}$ variation with $\mu_{S}$ in the case of $K=0$ in Fig. 3. As expected, the $\kappa_{2}^{B}$ does not vary with the $\mu_{S}$ for $G_{V}=0$, as there is no flavor interdependency from both $K$ and $G_{V}$. Whereas for a finite $G_{V}$, the introduction of a positive strangeness chemical potential $\mu_{S}$, increases the abundance of anti-strange quarks in the system for a fixed baryon chemical potential, $\mu_{B}$. In contrast, the number densities of up and down quarks remain positive as they are only dependent on $\mu_{B}$. As a result, increasing $\mu_{S}$ reduces the total number density, thereby diminishing the effect of the vector interaction [see Eq. (5)]. This leads to a cancellation of the vector interaction effect, $G_{V}$, at a specific value of $\mu_{S}$, resulting in this crossing.
It is important to note that the crossing points differ between Model-I and Model-II. In Model-II, this crossing corresponds to a negative value of $\kappa_{2}^{B}$ and occurs at a higher value of $\mu_{S}$. This suggests that $\mu_{S}$ has a weaker influence on the chiral transition line in the presence of flavor-dependent vector interactions $(g_{V}\neq 0)$. Moreover, for $K=0$, the flavor-dependent interaction related to $\mu_{S}$ does not impact the light quark chemical potential, leading to a curvature coefficient that remains independent of $\mu_{S}$. This is demonstrated by the two straight lines, representing $g_{V}^{+}$ (soft blue) and $g_{V}^{-}$ (soft orange), in Fig. 3. This indicates that in Model-II, there is no flavor mixing due to nonzero $g_{V}$ and the reduction in $\kappa_{2}^{B}$ with increasing $\mu_{S}$, as presented in the lower panel Fig.2, is entirely due to the flavor mixing term $K$.
III.3 $\mu_{S}$ variation of $\kappa_{4}^{B}$
Figure 4: The curvature coefficient for different values of $\mu_{S}$ in Model II with the same parameter sets used in the above graph.
Although, in scenarios considered in LQCD, the curvature coefficient $\kappa_{4}^{B}$ are obtained to be zero, we have obtained a significant dependence on $\mu_{S}$. Fig. 4 represents the models’ predictions on $\kappa_{4}^{B}$ as a function of $\mu_{S}$. The overall behaviour of $\kappa_{4}^{B}$ as a function of $\mu_{S}$ is the same (increasing) for both Model-I and Model-II with the corresponding attractive interactions $(G_{V}^{-}/g_{V}^{-})$ producing a higher curvature as compared to repulsive ones for a given value of $\mu_{S}$. This observation might provide us with a better understanding of the effect of $\mu_{S}$ as well as the vector interaction on the crossover line once explored in LQCD.
III.4 Effect of vector interaction on the critical endpoint
Figure 5: Effect of the vector interactions on the QCD phase line for $\mu_{S}=0$. Blue and magenta lines denote the estimation from the upper and lower bounds of the flavor-dependent vector interaction $g_{V}$, respectively. Soft blue and soft orange denote the same for flavor-independent scenario, i.e., with nonzero $G_{V}$. The red line denotes the limiting case with no vector interaction, $g_{V}=G_{V}=0$. The circles denote the CEP for the corresponding lines. Closed and open symbols are for Model-I and Model-II, respectively. (The line overlap for $G^{+}_{V}$ and $g^{+}_{V}$ at low $\mu_{B}$).
For physical light quark masses, the QCD transition is expected to be a crossover at zero or low baryon densities. However, at high baryon densities the transition is expected to be first order. Thus, the crossover is expected to meet the first order phase transition at some critical end point (CEP). It is important to investigate the impact of constrained vector interaction strengths on the location of this CEP. Previous studies have shown that the CEP’s position shifts based on the strength of the vector interaction [19, 22, 51]. Non-zero couplings, $G_{V}$, and $g_{V}$, both affect the curvature coefficients and the extension of the pseudo-critical (cross-over) region in the phase diagram.
In Fig. 5, we show the evaluated phase lines for constrained values of $G_{V}$ and $g_{V}$, with the corresponding CEPs indicated by circle symbols. The exact locations of the CEPs are provided in Table 4. As the repulsive vector interaction flattens the phase boundary, the CEP is expected to shift rightward for the positive upper limits of both $G_{V}$ and $g_{V}$, resulting in a corresponding decrease in $T^{\text{CEP}}$, as seen for the $g_{V}^{+}$ line. It is important to note that in Model-I, all three quark flavors contribute to the shifted chemical potential $\tilde{\mu}$. In contrast, Model-II includes flavor-dependent contributions, where only individual quark flavors affect $\tilde{\mu}$. As a result, the phase boundary is more compressed in the flavor-independent scenario. These results align with those in Ref. [19]. For the negative lower values of $G_{V}$ and $g_{V}$, the CEP shifts leftward, with an associated increase in $T^{\text{CEP}}$. It should be emphasized that the locations of these CEPs do not contradict predictions from Lattice QCD [7] or FRG-based studies [8], but instead offer a baseline within this effective model framework. Future constraints on the CEP location from first-principle QCD calculations may help narrow or exclude the range of coupling strengths explored here.
$G_{V},g_{V}$ $\mu^{CEP}_{B}$ (GeV) $T^{CEP}$ (GeV)
$G_{V}=g_{V}=0$ 0.972 0.0481
$G^{-}_{V}$ 0.891 0.0635
$G^{+}_{V}$ NA NA
$g^{-}_{V}$ 0.870 0.0679
$g^{+}_{V}$ NA NA
Table 4: Locations of the CEP for different values of $G_{V}$ and $g_{V}$.
IV Summary and Outlook
In this work, we investigate the effect of vector-type interactions within the $2+1$-flavor Nambu–Jona-Lasinio (NJL) model, considering both flavor-independent and flavor-dependent interactions with coefficients $G_{V}$ and $g_{V}$, respectively. These interactions arise only in the finite-density regime, making it difficult to fix the strengths of $G_{V}$ and $g_{V}$ from the QCD vacuum properties. Previous attempts to constrain these coefficients using LQCD-estimated curvature coefficients of the crossover line were limited due to large uncertainties in the lattice results. Here, we revisit this issue and utilize the precise lattice results for $\kappa_{2}^{B}$ for the cases $\mu_{S}=0$ and $n_{s}=0$ to constrain $G_{V}$ and $g_{V}$. We varied these coefficients in units of $G_{S}$ over a range of negative to positive values and evaluated $\kappa_{2,4}^{B}$ for the chiral crossover line. Negative values of $G_{V}$ and $g_{V}$ produce larger values of $\kappa_{2}^{B}$, which decrease as $G_{V}/G_{S}$ and $g_{V}/G_{S}$ increase towards positive values and further. This approach provides a tighter constraint on the range of these interaction coefficients. In the $\mu_{S}=0$ case, the allowed ranges of interaction strength are $\left[-0.117,0.212\right]$ and $\left[-0.296,0.552\right]$ (in units of $G_{S}$) for the flavor-independent and flavor-dependent interactions, respectively. Conversely, the $n_{s}=0$ case provides narrower ranges: $\left[-0.038,0.334\right]$ for $G_{V}/G_{S}$ and $\left[-0.076,0.667\right]$ for $g_{V}/G_{S}$. As the estimates for $\kappa_{4}^{B}$ are consistent with lattice results and remain within variances [4] for a wide range of $G_{V},g_{V}$, they do not further constrain the vector interaction strength. At this juncture, we want to iterate that the bound on $G_{V},g_{V}$ we get for the scenarios $\mu_{S}=0$ and $n_{s}=0$ are different, which necessitates the consideration of a more realistic interaction strength which depends on $\mu$.
We find that the improved lattice data on $\kappa_{2}^{B}$ provides narrower bounds on the vector interactions and can accommodate both attractive and repulsive vector interaction within the model framework. We test the model’s prediction in different physical scenarios, such as zero strangeness or the strangeness neutrality condition. $\kappa_{2}^{B}$ is always smaller in the latter case for both increasing $G_{V}$ and $g_{V}$, except at higher values of $G_{V}$ (flavor-independent scenario), where we observe a crossing. However, the bound is always wider on $g_{V}$ as compared to $G_{V}$, which can be explained within the model framework. The models’ predictions on the fourth-order curvature coefficient $(k_{4}^{B})$ match with the corresponding LQCD results, which can be termed as consistent with zero with the error bars.
In the NJL model, the mixing between light and strange quarks primarily arises from the ’t Hooft determinant term, introduced to take into account the QCD axial anomaly. The flavor-independent vector interaction also couples the light and strange quark sectors through their net densities. In this work, we have proposed a novel method to separate out the contribution of the flavor-independent vector interaction to flavor mixing by investigating the effect of the strangeness chemical potential, $\mu_{S}$, on $\kappa_{2}^{B}$, which was first proposed in Ref. [27]. These flavor mixings between light and strange sectors are responsible for the decrease in $\kappa_{2}^{B}$ as $\mu_{S}$ increases. Our findings are in general agreement with those from Ref. [27] We also observe that at a specific value of $\mu_{S}$, the effect of the vector interactions become negligible for both Model-I and Model-II and match with those for $G_{V}=g_{V}=0$ limit. This value of $\mu_{S}$ is larger for Model-II, signifying weak flavor mixing in the flavor-dependent coupling as compared to the flavor-independent one. Such interesting characteristics provide us with a better understanding of the types of vector interactions and possible signatures to distinguish them.
To assist the previous understanding, we have studied the case where the ’t Hooft determinant term is zero. This has enabled us to explicitly demonstrate the impact of vector interactions on the variation of $\kappa_{2}^{B}$ and allow for direct assessment of the flavor mixing driven by this flavor-independent vector interaction. In this limit, there is no variation of $\kappa_{2}^{B}$ in model II, indicating no mixing effect for the flavor-dependent vector interaction.
For the final part of our study, we find the location of the CEP for the allowed range of vector interactions. The location of CEP is heavily impacted by the strength of the vector interaction. We find that the CEP exists for the negative lower bounds of the vector interactions and disappears for the positive upper bounds in both the Model-I and Model-II. Although the model’s predictions on the location of CEP are consistent with the existing LQCD bound [7, 8], one should keep in mind that such observations can, in principle, depend on multiple factors, capturing all of which remains a task beyond such simple effective model calculations.
Acknowledgment
D.B. is supported by the Department of Science and Technology, Government of INDIA under the SERB National Post-Doctoral Fellowship Reference No. PDF/2023/001762. C.A.I. acknowledges Goethe University’s R3 Career Support, supported by Johanna Quandt Young Academy @@@@Goethe (JQYA) and the support by the State of Hesse within the Research Cluster ELEMENTS with project ID 500/10.006. He also thanks ECT* and CRC for support at the Workshop “Spin and quantum features of QCD plasma” during which this work has been completed.
Appendix A Variation of $\kappa_{4}^{B}$ with $G_{V}$ and $g_{V}$
Figure 6: The curvature coefficients, $\kappa_{4}^{B}$ as a function of $G_{V}/G_{S}$ (upper panel) $g_{V}/G_{S}$ (lower panel), respectively. The LQCD estimation of the same is $0.001(7)$ [4].
Appendix B $\kappa_{2}^{S}$ as a function of $G_{V}$ and $g_{V}$
Figure 7: The curvature coefficients, $\kappa_{2}^{S}$ as a function of $G_{V}/G_{S}$ (upper panel) $g_{V}/G_{S}$ (lower panel), respectively. The LQCD estimation of the same is $0.017(5)$ [4].